# Regularity and stability in a strongly degenerate nonlinear diffusion and haptotaxis model of cancer invasion

### v1

@@ 6 Discussion
We focused on a strongly degenerate nonlinear diffusion and haptotaxis model of cancer invasion, implementing the ECM-dependent saturation of the speed of cell motion due to the physical limits of cell migration. The resulting equation for the cell volume fraction dynamics exhibits two regimes of degeneracy of the velocity, which can make it particularly challenging to obtain sufficient regularity results to investigate the well-posedness of the system. We here devised specific regularity estimates, in finite time, that are compatible with these degeneracies. After some regularity results under physical bounds on the initial ECM volume fraction $\varphi_{0}$ (i.e. $0\leq\varphi_{0}\leq 1$), within which both regimes of degeneracy may be encountered, we obtained additional regularity estimates under the stricter bounds $0\leq\varphi_{0}\leq\frac{2}{3}$, focusing on the vacuum observed at $\varphi=0$. Based on these estimates and the associated local compactness results obtained, we proved the stability of – appropriately defined – weak solutions with respect to perturbations of the initial data. This confers robustness to the proposed cancer invasion model, despite the strong degeneracy. Moreover, the estimates derived in this work may also be useful to prove existence and uniqueness of solutions of the system, finalising the well-posedness of the model.
@@
Here, $\varphi_{min}>0$ is the minimum ECM volume fraction below which cell movement through the ECM is hindered, and $\varphi_{th}>\varphi_{0}$ is a threshold ECM volume fraction above which cells can no longer pass through the ECM pores. Definition (3) consists in the specific case in which $\varphi_{min}=0$ and $\varphi_{th}=1$, but different values will be more biologically relevant, in which case the degeneracy will occur for intermediate values of $\varphi$.

### v2

@@ 7 Discussion
We focused on a strongly degenerate nonlinear diffusion and haptotaxis model of cancer invasion, implementing the ECM-dependent saturation of the speed of cell motion due to the physical limits of cell migration. The resulting equation for the cell volume fraction dynamics exhibits two regimes of degeneracy of the velocity, which can make it particularly challenging to obtain sufficient regularity results to investigate the well-posedness of the system. We here devised specific regularity estimates, in finite time, that are compatible with these degeneracies. After some regularity results under physical bounds on the initial ECM volume fraction $\varphi_{0}$ (i.e. $0\leq\varphi_{0}\leq 1$), within which both regimes of degeneracy may be encountered, we obtained additional regularity estimates under the stricter bounds $0\leq\varphi_{0}\leq\frac{2}{3}$, focusing on the vacuum observed at $\varphi=0$. Based on these estimates and the associated local compactness results obtained, we proved the stability of – appropriately defined – weak solutions with respect to perturbations of the initial data. This confers robustness to the proposed cancer invasion model, despite the strong degeneracy. Finally, global existence of weak solutions was proved from an approximate non-degenerate problem, exploiting simple extensions of the preceding regularity results. Uniqueness and large-time behaviour of solutions of the strongly degenerate system are open problems.
@@
Here, $\varphi_{min}>0$ is the minimum ECM volume fraction below which cell movement through the ECM is hindered, and $\varphi_{th}>\varphi_{0}$ is a threshold ECM volume fraction above which cells can no longer pass through the ECM pores. Definition (3) consists in the specific case in which $\varphi_{min}=0$ and $\varphi_{th}=1$, but different values will be more biologically relevant, in which case the degeneracy will occur for intermediate values of $\varphi$ – see also Figure 2.
Figure 2: Alternative behaviours of the function $S(\varphi)$ based on definition (7) proposed in [15, 22], for $\varphi_{th}>\frac{2}{3}$ (left) or $\varphi_{th}<\frac{2}{3}$ (right), displaying wider regions of strong degeneracy (in red) of the velocity $\bm{V}$ defined in (2).
Finally, we remark that discretised versions of the estimates obtained in this work will be exploited to design and analyse a robust numerical scheme based on finite volume approximations to simulate the system. The scheme should effectively address the stiffness of the problem by preserving non-negativity, avoiding spurious oscillations, and accurately capturing the sharp gradients of solutions typical of cell migration PDE models [9, 19, 21], particularly critical for nonlinear and degenerate parabolic equations [12, 20], meanwhile conserving the qualitative properties of the system in the degenerate regimes [6, 7].